# LLM-Guided Open RAN: Empowering Hierarchical RAN Intelligent Control

### v1

@@ Abstract
Recent advancements in large language models have led to a significant interest in deploying LLM-empowered algorithms for wireless communication networks. Meanwhile, open radio access network (O-RAN) techniques offer unprecedented flexibility, with the (non-RT RIC) and (near-RT RIC) components enabling intelligent resource management across different time scales. In this paper, we propose the LLM empowered hierarchical RIC (LLM-hRIC) framework to improve the collaboration between RICs. This framework integrates LLMs with reinforcement learning (RL) for efficient network resource management. In this framework, LLMs-empowered non-RT RICs provide strategic guidance and high-level policies based on environmental context. Concurrently, RL-empowered near-RT RICs perform low-latency tasks based on strategic guidance and local near-RT observation. We evaluate the LLM-hRIC framework in an integrated access and backhaul (IAB) network setting. Simulation results demonstrate that the proposed framework achieves superior performance. Finally, we discuss the key future challenges in applying LLMs to O-RAN.
@@
Open radio access network, large language model, reinforcement learning
@@ I Introduction
A core component of O-RAN is the RICs, which provides advanced control and optimization capabilities. The RICs include: 1) the non-RT RIC, which integrates with the network orchestrator and operates on a time scale larger than 1 s, and 2) the near-RT RIC, which manages control loops with RAN nodes on a time scale between $10$ ms and $1$ s. Those RICs are capable of aggregating key performance measurements (KPMs) to monitor the status of network infrastructure (e.g., number of users, load, throughput, resource utilization). Specifically, the near-RT RIC connects to the RAN components through E2 interface. The non-RT RIC interfaces with RAN components through O1 interface, and based on the available KPMs, it facilitates a non-real-time control loop and supports the deployment of policies, guidance, and intelligent models within the near-RT RIC via A1 interface. The third-party applications, known as xApp and rApp, are deployed within near-RT RIC and non-RT RIC, respectively. These applications enable advanced functionalities, such as enhanced radio resource management, improved energy efficiency, and predictive maintenance [2].
@@
Significant researches have been conducted to enhance RICs functionalities [3, 4, 5, 6, 7, 8]. Specifically, near-RT RIC, equipped with xApps, manages a smaller number of devices and enable low-latency optimizations. These optimizations include radio card switching to improve energy efficiency [3], flow-split distribution, congestion control, and scheduling to optimize network utility [4]. In contrast, the non-RT RIC, equipped with rApps, is responsible for managing a larger number of devices and performing high-level, long-term network management [5, 6, 7, 8]. For example, the orchestRAN framework in [5] automatically computes the optimal set of data-driven algorithms and determines their execution locations. Furthermore, cooperation between non-RT RIC and near-RT RIC is investigated in [6, 7, 8]. Specifically, the non-RT RIC provides near-RT RIC with high-level guidance, including network configuration and reconfiguration information [6], control signals [7], and optimized control functions [8]. The near-RT RIC then leverages real-time information to refine and implement these policies, enhancing dynamic network performance. Despite these advancements, two critical challenges remain: 1) Traditional machine learning (ML) techniques struggle to process the diverse and complex information in O-RAN. 2) Ensuring high-quality policy generation in the non-RT RIC across diverse scenarios and improving efficient cooperation between non-RT RIC and near-RT RIC are challenging.
Meanwhile, the LLMs, such as OpenAI’s ChatGPT, Google’s Gemini, and Meta AI’s Llama, are advanced ML models, designed to understand and generate human language. Trained on vast and diverse datasets, LLMs excel in tasks such as text generation, translation, summarization, and question answering. Their ability to understand context and generate contextually relevant outputs makes them highly adaptable across various domains. In the context of telecommunications, LLMs offer unique advantages by processing and analyzing extensive network data to improve decision-making. Telecom operators can leverage LLMs to enhance network management and improve overall service quality. Additionally, LLMs support the development of intelligent applications for network operations, including predictive maintenance, anomaly detection, and performance optimization. The LLMs have opened new routes for wireless network design and operation, offering the potential for significant performance improvements.
Recent studies have applied LLM-based technologies to wireless network for resource allocation[9, 10, 11], network deployment[12, 13] and packet analysis [14]. Specifically, [9] develope a LLM empowered-xApp to allocate physical resource block (PRB) for 5G O-RAN. However it lacks a comparison with state of art resource allocation algorithms and does not analyze the inference latency of LLMs. [10] and [11] utilize Meta AI’s Llama to derive the power allocation policy directly. Additionally, [12] deploys a combined LLM and conventional neural network (CNN) model as a RL agent for base station placement. Furthermore, the GPT model has been utilized to formulate mathematical models and generate solution codes for base station placement [13]. Although these applications show the potential of LLMs in wireless networks, several challenges remain. First, the high computational complexity hinders near-RT decision-making. Additionally, without fine-tuning for domain-specific tasks, their performance often remains comparable to traditional algorithms.
Motivated by these opportunities and challenges, we propose the LLM-hRIC framework for O-RAN. This framework deploys LLMs-empowered non-RT RIC to provide guidance for the RL-empowered near-RT RIC. The near-RT RIC utilizes this guidance along with near-RT environment information to perform near-RT decision-making. The contributions of this work are summarized as follows:
  • LLM-empowered non-RT RIC: We propose a LLM-empowered non-RT RIC that collects and processes comprehensive environmental data to provide strategic guidance to the near-RT RIC. This approach improves data integration and long-term decision-making in the O-RAN framework.
  • RL-empowered near-RT RIC: We introduce a RL-empowered near-RT RIC that utilizes the guidance from non-RT RIC and local observations to make near-RT decisions. To accelerate convergence and improve overall system performance, we deploy a joint training method that smoothly mixes non-RT RIC’s guidance and near-RT RIC’s policy.
  • Performance Evaluation: We evaluate the proposed framework in IAB networks. The results show significant performance improvements compared to baseline methods, demonstrating its ability to manage the complex information in O-RAN systems.
@@ II LLM-empowerd hierarchical RIC Framework
In this section, we propose the LLM-empowerd hierarchical RIC (LLM-hRIC) framework for the resource management in O-RAN as shown in Fig. 1. The LLM-hRIC framework consists of 1) the LLM-empowered non-RT RIC (i.e., upper layer control) and 2) the RL-empowered near-RT RIC (i.e., lower layer control), and this shows how the LLM can be used as a way to facilitate the resource management in O-RAN. As the LLM has the advantages on generalization across various environments with minimal language data, it is used for the non-RT RIC to provide a guidance on the strategies to the near-RT RIC from the large-scale view of the networks. The near-RT RIC can be designed to utilize this guidance in conjunction with the small-scale view (i.e., local information of RAN) to make decisions in near-RT. More details on the LLM-hRIC framework are provided in the following subsections.
Figure 1: Proposed LLM-hRIC framework.
@@ II-A LLM-Empowered non-RT RIC
The non-RT RIC serves as a strategic layer, which works based on the collected KPMs from various network units including gNB and near-RT RIC via O1 interface. The rApps, deployed for the non-RT RIC, is responsible for utilizing those information to generate strategic guidance for the near-RT RIC. By integrating LLM technology, the non-RT RIC gains the capability to process and analyze multimodal data including text, images, and signals, enables to form a comprehensive understanding of the current network environments. Additionally, LLMs excels in synthesizing diverse data sources and leveraging their extensive parameters and deep architecture to perform complex reasoning and inference. These strengths allow the LLM-empowered non-RT RIC to tackle challenges in complex network management and provide a high-quality strategic guidance to the near-RT RIC.
@@
    Data Integration: The non-RT RIC collects multimodal data (e.g., textual report on network conditions and radio spectrum images), and integrates the collected data by structuring it into an unified format.
@@
    Environment Understanding and Actionable Insights Generation: The LLM analyze and understand the context based on the integrated multimodal data, and generate a strategic guidance (e.g., transmit power allocation strategies), which can be helpful to make decisions in the near-RT RIC (e.g., RL-based power control). The mathematical formulation of the guidance generation process is given as
@@
    where $x_{n}\in\mathcal{X}$ represents a specific data modality in the input set $\mathcal{X}$, and $\theta$ denotes the parameters of LLM.
@@
    Guidance transfer to the near-RT RIC: The guidance vector y is then sent to the near-RT RIC via A1 interface to be used for the near-RT decision-making process.
@@ II-B RL-Empowered near-RT RIC
The near-RT RIC operates as a politic layer of O-RAN, managing and optimizing network operations in near-RT. To achieve this, RL-based xApps can be deployed for near-RT RIC to determine optimal actions based on 1) the observed local information $\boldsymbol{o}$ via E2 interface and 2) the guidance vector y, received from the LLM-empowered non-RT RIC via A1 interface. Specifically, the optimal action $\boldsymbol{a}$ can be determined as
@@
Generally, given the strict latency constraints of the near-RT RIC (e.g., 10 ms to 1s), the RL-based xApps need to be designed to process information efficiently and make a decision fast. From the integration of LLM-empowered non-RT RIC, the RL component’s ability to refine its decision-making can be enhanced due to the LLM-generated insights (i.e., guidances), obtained from the information on large-scale network environments. This can enable the RL-based xApps to make a decision with the large-scale view on the network (e.g., overall traffic distribution and inter-cell interference) as well as the localized and real-time observation (e.g., instantaneous user demand and channel quality) of the network.
@@ III Use Case: Power Allocation in IAB Network
III-A Power Allocation for IAB Networks
The IAB networks as illustrated in Fig. 1, consists of a set of macro base stations, denoted as =$\mathcal{M}=\{1,\cdots,M\}$, distributed in a network. Each MBS is connected with $N$ small base stations, and $\mathcal{N}_{m}=\{1,\cdots,N\}$ denotes the set of SBSs, connected to the $m$-th MBS. A set of mobile users, $\mathcal{K}$, follow a Gaussian-Markov mobility model and always connect to the nearest SBS. The SBS allocates fixed power to their connected user.
The interference exists among SBSs connected to different MBS that use the same sub-carrier. Both line-of-sight (LoS) and non-line-of-sight (NLoS) path loss characteristics are considered, where the probability of a LoS connection is modeled as $P^{\text{Los}}=\text{exp}(-d/\rho)$, where $d$ represents the distance between MBS and SBS, and $\rho$ is a LoS range constant. The large-scale fading gain of LoS and NLoS components are given by $L^{\text{LoS}}$ and $L^{\text{NLoS}}$, respectively. Furthermore, the small-scale fading is modeled as independent Nakagami-$m$ fading.
@@ III-B1 LLM-Empowered non-RT RIC
  • Constraints: Constraints for power allocation policy and the guidance format.
The structure of the prompt is as follows:
@@
Due to the time-varying nature of small-scale fading, the non-RT RIC uses average values of small-scale fading parameters to compute metrics like channel gain, expected data rate, and interference. These average values effectively represent the channel’s state in future time slots. The LLM generates the initial power allocation policy as a guidance, denoted as $\boldsymbol{p}^{o}_{m}$, for the $m$-th MBS. This guidance is updated every $T_{0}$ time slots to adapt to network environment change.
@@ III-B2 RL-Empowered near-RT RIC
  • Reward Function: The reward is designed to encourage individual performance and global cooperation $r_{m}=r^{\text{l}}_{m}+r^{\text{g}}$, where $r^{\text{l}}_{m}$ is the throughput of $m$-th MBS, and $r^{\text{g}}$ is the total throughput of all MBSs.
III-C Cooperative Training
The RL agent in the near-RT RIC need to be trained to effectively utilize the guidance provide by LLM. Traditional RL typically relies on random noise for action exploration during the initial training stages. To enhance exploration efficiency, the initial actions are instead guided by the LLM, e.g., initial power allocation policy. And then the RL action will mix with LLM guidance gradually. The training is divided into the following three phases:
  • Phase 1: The power allocation policy $\boldsymbol{p}_{m}$ is generated by adding noise to the policy $\boldsymbol{p}_{m}^{o}$ provided by the LLM, facilitating exploration around the LLM’s recommendations.
  • Phase 2: The power allocation policy $\boldsymbol{p}_{m}$ is obtained by linearly blending the LLM-provided policy $\boldsymbol{p}_{m}^{o}$ and the RL-generated policy $\boldsymbol{p}_{m}^{d}$ using a decay weighted $w$: $\boldsymbol{p}_{m}=w\boldsymbol{p}_{m}^{o}+(1-w)\boldsymbol{p}_{m}^{d}$, where $w$ gradually decays from $1$ to 00 during this phase.
  • Phase 3: The power allocation policy $\boldsymbol{p}_{m}$ is determined by the RL agent.
In each training stage, the power allocation policy interact with environment, and the resulting MDP experiences are collected to continuously train the RL agent.
@@ III-D Experimental Results
In this simulation, Llama-3.1-8B-instruct model [15] is deployed in the rApp to generate a power allocation guidance policy in a large time scale (e.g., 1s). Meanwhile, a deep deterministic policy gradient (DDPG) agent is deployed in the xApp to provide fine-grained power allocation policy for a MBS in a small time scale (e.g., 100ms). To assess the effectiveness of the proposed framework, we compare it against the following baseline methods: 1) DDPG with linear decay noise for action exploration (DLN), 2) DDPG with cosine decay noise for action exploration (DCN), and 3) equal power allocation (EPA).
Figure 2: Training curve for all methods when $\alpha=0.5$.
@@
Figure 3: Effect of backhaul-access bandwidth partitioning ratio $\alpha$ on total throughput.

### v2

# LLM-hRIC: LLM-empowered Hierarchical RAN Intelligent Control for O-RAN

@@ Abstract
Despite recent advances in applying large language models and machine learning (ML) techniques to open radio access network (O-RAN), critical challenges remain, such as insufficient cooperation between RICs, high computational demands hindering real-time decisions, and the lack of domain-specific fine-tuning. Therefore, this article introduces the LLM-empowered hierarchical RIC (LLM-hRIC) framework to improve the collaboration between RICs in O-RAN. The LLM-empowered non-real-time (non-RT RIC) acts as a guider, offering a strategic guidance to the near-real-time (near-RT RIC) using global network information. The RL-empowered near-RT RIC acts as an implementer, combining this guidance with local real-time data to make near-RT decisions. We evaluate the feasibility and performance of the LLM-hRIC framework in an integrated access and backhaul (IAB) network setting, and finally, discuss the open challenges of the LLM-hRIC framework for O-RAN.
@@
Open radio access network, large language model, RAN intelligent control, reinforcement learning
@@ I Introduction
A core component of O-RAN is the RICs, which provides advanced control and optimization capabilities. The RICs include: 1) the non-RT RIC, which integrates with the network orchestrator and operates on a time scale larger than 1 s, and 2) the near-RT RIC, which manages control loops with RAN nodes on a time scale between $10$ ms and $1$ s. The non-RT RIC and the near-RT RIC are capable of obtaining key performance measurements (KPMs) via O1 and E2 interfaces, respectively, to monitor the status of network infrastructure (e.g., number of users, load, throughput, and resource utilization). The non-RT RIC also supports the deployment of policies, guidance, and intelligent models within the near-RT RIC via the A1 interface. Specific RIC functionalities, such as mobility management and network slicing, are deployed within near-RT RIC and non-RT RIC in the form of the third-party applications, known as xApp and rApp, respectively [2].
@@
Recently, intensive researches on RICs have been appeared[3, 4, 5, 6, 7, 8]. Specifically, the near-RT RICs has been developed to manage a smaller number of devices and optimize on a small-time scale, such as the radio card switching for energy efficiency [3] and the flow-split distribution, congestion control, and scheduling for higher network utility [4]. The non-RT RIC has been developed to manage a larger number of devices and perform high-level and long-term network management, such as the orchestRAN[5], which automatically computes the optimal set of algorithms and their execution location. Furthermore, the cooperation between non-RT RIC and near-RT RIC is investigated in [6, 7, 8]. Specifically, the non-RT RIC provides the high-level guidance to near-RT RIC, including the network configuration and reconfiguration information [6], control signals [7], and optimized control functions [8]. The near-RT RIC then refines and implements these policies based on real-time information to enhance the network performance.
Meanwhile, the LLMs have opened new routes for wireless network design and operation including O-RAN. The LLMs are advanced ML models, designed to understand and generate human language. Their ability to understand context and generate contextually relevant outputs makes them highly adaptable across various domains. In the context of telecommunications, LLMs offer unique advantages by processing and analyzing extensive network data to improve decision-making. Recent studies have applied LLM-based technologies to wireless network for the resource allocation[9, 10], base station placement[11, 12], and packet analysis [13]. Recently, the LLM-based xApp to allocate physical resource block (PRB) for O-RAN has also been introduced[14].
Despite these advancements in RICs and LLM-based techniques for wireless networks, critical challenges still remain as follows: 1) traditional ML techniques struggle to process the diverse and complex information in O-RAN; 2) the efficient cooperation between non-RT RIC and near-RT RIC still remains largely unexplored; 3) although the potential on the role of LLMs in wireless networks has been shown, the high computational complexity hinders near-RT decision-making; and 4) without fine-tuning for domain-specific tasks, the performance of LLM-based techniques often remains comparable to or worse than traditional algorithms.
Motivated by these opportunities and challenges, we propose the novel framework for O-RAN, which is the LLM-hRIC. In this framework, there are two RICs: 1) the LLMs-empowered non-RT RIC and 2) the reinforcement learning (RL)-empowered near-RT RIC. The LLMs-empowered non-RT RIC works as a guider to the near-RT RIC. Based on the objective of the near-RT RIC, it provides strategic guidance for the RL-empowered near-RT RIC based on the global view of the networks (e.g., long-term channel information, network nodes state, users mobility model, and users distribution). The RL-empowered near-RT RIC works as an implementer, which utilizes this guidance along with the local view of the networks (e.g., real-time channel quality and local network node state) to perform near-RT decision-making. The key contribution of the paper can be summarized as follows.
  • We propose the novel LLM-hRIC framework for O-RAN, which can show how the LLM can be utilized to enhance the performance of RICs in O-RAN. This article focuses on addressing the limitation of LLM in making near-RT decisions and paves a way to utilize it in a reliable way in O-RAN environments.
  • We introduce the three sequential phases for training of the RL-empowered near-RT RIC, of each phase utilizes the guidance of the non-RT RIC in a different way to accelerate convergence as well as improve the system performance.
  • We quantify the performance of the LLM-hRIC for the power allocation in the IAB networks as a use case. This can demonstrate the ability of LLM-hRIC to manage the complex information in O-RAN.
@@ II LLM-empowerd hierarchical RIC Framework
In this section, we introduce the LLM-empowered hierarchical RIC (LLM-hRIC) framework for resource management in O-RAN as shown in Fig. 1. The LLM-hRIC framework consists of 1) the LLM-empowered non-RT RIC (i.e., upper layer control) and 2) the RL-empowered near-RT RIC (i.e., lower layer control), and this shows how the LLM can be used as a way to facilitate the resource management in O-RAN. As the LLM has the advantages of generalization across various environments with minimal language data, it is used for the non-RT RIC to provide guidance on the strategies to the near-RT RIC from the large-scale view of the networks. The near-RT RIC can be designed to utilize this guidance in conjunction with the small-scale view (i.e., local information of RAN) to make decisions in near-RT. More details on the LLM-hRIC framework are provided in the following subsections.
Figure 1: Structure of the proposed LLM-hRIC framework.
@@ II-A LLM-Empowered non-RT RIC
The non-RT RIC serves as a strategic layer, which works based on the collected KPMs from various network units including gNB and near-RT RIC via O1 interface. The rApps, deployed for the non-RT RIC, are responsible for utilizing that information to generate strategic guidance for the near-RT RIC. By integrating LLM technology, the non-RT RIC gains the capability to process and analyze multimodal data including text, images, and signals, enabling to form of a comprehensive understanding of the current network environments. Additionally, LLMs excels in synthesizing diverse data sources and leveraging their extensive parameters and deep architecture to perform complex reasoning and inference. These strengths allow the LLM-empowered non-RT RIC to tackle challenges in complex network management and provide high-quality strategic guidance to the near-RT RIC.
@@
    Data Integration: The non-RT RIC collects multimodal data (e.g., textual report on network conditions and radio spectrum images) and integrates the collected data by structuring it into a unified format.
@@
    Environment Understanding and Actionable Insights Generation: The LLM analyzes and understands the context based on the integrated multimodal data, and generate strategic guidance (e.g., transmit power allocation strategies), which can be helpful to make decisions in the near-RT RIC (e.g., RL-based power control). The mathematical formulation of the guidance generation process is given as
@@
    where $x_{n}\in\mathcal{X}$ represents a specific data modality in the input set $\mathcal{X}$ and $\theta$ denotes the parameters of LLM.
@@
    Guidance Transfer to the near-RT RIC: The guidance vector y is then sent to the near-RT RIC via the A1 interface, used for the near-RT decision-making process.
@@ II-B RL-Empowered near-RT RIC
The near-RT RIC operates as a political layer of O-RAN, managing and optimizing network operations in near-RT. To achieve this, RL-based xApps can be deployed for near-RT RIC to determine optimal actions based on 1) the observed local information $\boldsymbol{o}$ via E2 interface and 2) the guidance vector y, received from the LLM-empowered non-RT RIC via A1 interface. Specifically, the optimal action $\boldsymbol{a}$ can be determined as
@@
Generally, given the strict latency constraints of the near-RT RIC (e.g., 10 ms to 1s), the RL-based xApps need to be designed to process information efficiently and make a decision fast. From the integration of LLM-empowered non-RT RIC, the RL component’s ability to refine its decision-making can be enhanced due to the LLM-generated insights (i.e., guidances), obtained from the information on large-scale network environments. This can enable the RL-based xApps to decide with a large-scale view of the network (e.g., overall traffic distribution and inter-cell interference) as well as the localized and real-time observation (e.g., instantaneous user demand and channel quality) of the network.
@@ III Use Case: Power Allocation in IAB Network
III-A Power Allocation in IAB Networks
The IAB networks as illustrated in Fig. 1, consists of a set of macro base stations, denoted as =$\mathcal{M}=\{1,\cdots,M\}$, distributed in a network. Each MBS is connected with $N$ small base stations, and $\mathcal{N}_{m}=\{1,\cdots,N\}$ denotes the set of SBSs, connected to the $m$-th MBS. A set of mobile users $\mathcal{K}_{m,n}$, connected with the $n$-th SBS of the $m$-th MBS, follow a Gaussian-Markov mobility model. The SBS allocates fixed power to their connected user. The interference exists among SBSs connected to different MBS that use the same sub-carrier. Both line-of-sight (LoS) and non-line-of-sight (NLoS) path loss characteristics are considered, where the probability of a LoS connection is modeled as $P^{\text{Los}}=\text{exp}(-d/\rho)$, where $d$ represents the distance between MBS and SBS, and $\rho$ is a LoS range constant. Furthermore, the small-scale fading is modeled as independent Nakagami-$m$ fading.
@@ III-B1 LLM-Empowered non-RT RIC
  • Constraints: Details the constraints on power allocation policy and the guidance format.
The structure of the prompt is as Fig. 2.
@@
Figure 2: The Prompt Structure.
Due to the time-varying nature of small-scale fading, the non-RT RIC uses average values of small-scale fading parameters to compute metrics like channel gain, expected data rate, and interference. These average values effectively represent the channel’s state in future time slots.
@@ III-B2 RL-Empowered near-RT RIC
  • Reward Function: The reward is designed to encourage individual performance and global cooperation $r_{m}=r^{\text{l}}_{m}+r^{\text{g}}$, where $r^{\text{l}}_{m}$ is the throughput of $m$-th MBS and $r^{\text{g}}$ is the total throughput of all MBSs.
III-C LLM-assisted Three Phases Training for RL
The RL agent in the near-RT RIC must be trained to effectively utilize the guidance provided by LLM. Traditional RL typically relies on a random policy for action exploration during the initial stage of training. To improve the exploration efficiency, this random policy is replaced with the LLM guidance, such as an initial power allocation policy and injected action noise. The policy is gradually formed by blending the actions, generated by the RL agent with the LLM guidance, until the agent is capable of making decisions independently. Specifically, the training can be processed by the following three sequential phases:
  • Phase 1 (LLM-guided Exploration): The power allocation policy $\boldsymbol{p}_{m}$ is determined by adding noise to the LLM’s initial guidance policy $\boldsymbol{p}_{m}^{o}$, i.e., $\boldsymbol{p}_{m}=\boldsymbol{p}_{m}^{o}+\text{noise}$, to encourage exploration in the vicinity of the LLM’s guidance.
  • Phase 2 (Policy Blending): The power allocation policy $\boldsymbol{p}_{m}$ is determined by a weighted combination of the LLM-provided policy $\boldsymbol{p}_{m}^{o}$ and the RL-generated policy $\boldsymbol{p}_{m}^{d}$, i.e., $\boldsymbol{p}_{m}=w\boldsymbol{p}_{m}^{o}+(1-w)\boldsymbol{p}_{m}^{d}$, where the blending coefficient $w$ decays gradually from $1$ to 00 over time.
  • Phase 3 (Self-directed Decision): The power allocation policy $\boldsymbol{p}_{m}$ is determined solely by the RL agent, i.e., $\boldsymbol{p}_{m}=\boldsymbol{p}_{m}^{d}$.
In each training stage, the power allocation policy interacts with environment, and the resulting MDP experiences are collected to train the RL agent.
@@ III-D Experimental Results
In this simulation, Llama-3.1-8B-instruct model [15] is deployed in the rApp to generate a power allocation guidance policy in a large time scale (e.g., 2s). Meanwhile, a deep deterministic policy gradient (DDPG) agent is deployed in the xApp to provide fine-grained power allocation policy for a MBS in a small time scale (e.g., 200ms). To assess the effectiveness of the proposed framework, we compare it against the following baseline methods: 1) DDPG with linear decay noise for action exploration (DLN), 2) DDPG with cosine decay noise for action exploration (DCN), and 3) equal power allocation (EPA).
Figure 3: Training curve for proposed and baseline methods over 500 training epochs when $M=3,N=6,W=100\text{Mb},P_{m}^{\text{max}}=44\text{dBm}$, and $\alpha=0.5$.
@@
Figure 4: Average total throughput of the proposed LLM-hRIC and baseline methods according to backhaul-access bandwidth partitioning ratio $\alpha$ over 50 test examples.